DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, November 18 Padelford C-401 2:30 pm NOTE: Special day, time and room

The resolvent of the Laplacian on manifolds with asymptotically

conic ends

Andrew Hassell (Australian National University, Canberra)

An asymptotically conic manifold is a complete manifold with one end, which has a metric asymptotic to dr^{2} + r^{2}h as r tends to infinity. Here h is a metric on the end, assumed to be a compact manifold. Such a metric is known as a scattering metric. I will describe a direct construction of the resolvent, R(σ), of the Laplacian of a scattering metric, for σ on the positive real axis (ie, on the spectrum). This is joint work with Andras Vasy.